2, It is known that the quantitative relationship between the dependent variable Y and the independent variables X and 2 is: A. Make a Table and draw the graph of the relationship between Y and X when the numerical value of Z is: B. Make a Table and draw the graph of the relationship between Y and Z when the numerical value of X is: C. Make a Table of the value of the rate of change of Y with respect to change of X for values of x between negative five and positive five. 3. It is known that the relationship between the dependent variable Y and the independent variables X and Z is A. Make a Table and draw a graph of the relationship between Y and z when the value of X is: b.X1
Answer #1

2) Given that,

Y = X^2 + (2.Z)

A) For Z = 0,

Y = X^2

Table will be as mentioned below-

X Y
0 0
1 1
2 4
3 9
4 16
5 25
6 36
7 49

Graph will be as mentioned below-

For Z = -1

Y = X^2 - 2

Table will be as mentioned below -

X Y
0 -2
1 -1
2 2
3 7
4 14
5 23
6 34
7 47

Graph will be as mentioned below -

Abhishek Rai _ Chart Tools Book4 - Excel Formulas Data Review View Help Design Format Tell me what you want to do Home nsert Page Layout Gridlines Headings Colors A Fonts Width:Automatic Height: Automatic-| | View | 1x1 View Bring Send Selection Align Group Rotate Margins Orientation Size Print Breaks Background Print Area Print Print Forward Backward Pane Titles | Scale: 110% Effects Scale to Fit Sheet Options Page Setup Chart 1 12 13 15 17 20 21 Shee + 100% Display Settings 圈1回凹ー Average: 9.5 Count 18 Sum: 152 ENG US 5:38 PM 23-Jan-19 ^ :r x Type here to search

B)

For X =0,

Y = 2*Z

Table will be as mentioned below -

Z Y
0 0
1 2
2 4
3 6
4 8
5 10
6 12
7 14

Graph will be as mentioned below -

Book4 - Excel Chart Tools Abhishek Rai困- Home nsert Page Layout Formulas Data Review View Help Design Format Tell me what you want to do Width:Automatic Height: | Automatic-| | View | 1x1 View Colors Gridlines Headings A Fonts Margins Orientation Size Print Breaks Background Print Bring Send Selection Align Group Rotate □ Print | ■ print | Forward-Backward Effects Titles | Scale |100% Page Setup Scale to Fit Sheet Options Chart 1 16 12 12 17 20 Sheet1 | ④ Average: 5.25 Display Settings E |回凹ー Select destination and press ENTER or choose Paste Count: 18 Sum: 84 + 100% Type here to search ENG 5:42 PM US 23-Jan-19

For X = -5

Y = 25 + 2*Z

Table will be as mentioned below -

Z Y
0 25
1 27
2 29
3 31
4 33
5 35
6 37
7 39

Graph will be as mentioned below -

Book4 Excel Chart Tools Abhishek Rai -0 Home nsert Page Layout Formulas Data Review View Help Design Format Tell me what you want to do Colors Width:Automatic Gridlines Headings Height: | Automatic-| | View | 1x1 View A Fonts Automati Themes Effects Margins Orientation Size Print Breaks Background Print Area reaks Background Print □ Print | ■ print | Forward-Backward Bring Send Selection Align Group Rotate Titles | Scale |100% Pane Themes Page Setup Scale to Fit Sheet Options Chart 1 Y=25+2*z 13 17 21 Sheet1 | ④ Display Settings 回凹ー + 100% Type here to search :r ^ x US 23-Jan-19 45 40 35 30 25 20 15 10 5

C)

Suppose that Z has value equal to 0. Table will be as mentioned below -

X Y Change %
-5 25 100%
-4 16 64%
-3 9 36%
-2 4 16%
-1 1 4%
0 0 0%
1 1 4%
2 4 16%
3 9 36%
4 16 64%
5 25 100%